### v2

# IR2: Information Regularization for Information Retrieval thanks: Accepted by LREC-COLING 2024

Abstract
Effective information retrieval (IR) in settings with limited training data, particularly for complex queries, remains a challenging task. This paper introduces IR2, Information Regularization for Information Retrieval, a technique for reducing overfitting during synthetic data generation. This approach, representing a novel application of regularization techniques in synthetic data creation for IR, is tested on three recent IR tasks characterized by complex queries: DORIS-MAE, ArguAna, and WhatsThatBook. Experimental results indicate that our regularization techniques not only outperform previous synthetic query generation methods on the tasks considered but also reduce cost by up to 50%. Furthermore, this paper categorizes and explores three regularization methods at different stages of the query synthesis pipeline—input, prompt, and output—each offering varying degrees of performance improvement compared to models where no regularization is applied. This provides a systematic approach for optimizing synthetic data generation in data-limited, complex-query IR scenarios. All code, prompts and synthetic data are available at https://github.com/kwang927/Information-Regularization.
IR2: Information Regularization for Information Retrieval thanks: Accepted by LREC-COLING 2024
Jianyou Wang thanks: Equal contribution  Kaicheng Wang22footnotemark: 2  Xiaoyue Wang22footnotemark: 2  Weili Cao Ramamohan Paturi thanks: Equal Senior Authors  Leon Bergen33footnotemark: 3 Laboratory for Emerging Intelligence University of California, San Diego {jiw101, kaw036, xiw027, w2cao, rpaturi, lbergen}@ucsd.edu
1 Introduction
Users often submit complex queries to information retrieval (IR) systems, expecting answers that accurately address multiple, interconnected questions. These complex queries may involve ambiguous language or multiple specific criteria, requiring that IR systems decipher and respond to the layered intents accurately to provide relevant results Wang et al. (2023); Lin et al. (2023). Ensuring that IR systems effectively handle the subtleties of such queries is important, allowing these systems to align with varied and nuanced user demands.
Effectively addressing complex queries in IR depends on a model’s exposure to a wide range of user queries during training. However, obtaining diverse real-world training data is often constrained by privacy concerns, availability, and resources. Synthetic data, therefore, becomes crucial, offering a means to expand training datasets, enabling models to learn from a broader spectrum of queries and user intents Ma et al. (2021); Liang et al. (2020).
Recent research exploits Large Language Models (LLMs) to generate synthetic data pairs, constructing synthetic queries from real passages, often derived from zero-shot or few-shot examples Bonifacio et al. (2022); Jeronymo et al. (2023); Meng et al. (2022); Peng et al. (2023); Penha et al. (2023). Addressing the challenges of complex query information retrieval (IR) tasks through LLM-based synthetic data generation presents distinct difficulties. While synthetic data improves model performance across multiple tasks and metrics, generating queries from documents often results in synthetic pairs that exhibit superficial textual patterns, such as synonyms, matching keywords, and similar organizational flow, see the Promptagator Query in Figure 2 by a standard 8-shot synthetic query generation technique Dai et al. (2022). Though this might be effective for certain IR tasks, models may overfit to more superficial features, preventing them from understanding more conceptual relationships between query and document.
Figure 1: Performance of synthetic data generation methods on complex IR benchmarks. The $\triangle$, $\square$ Gao et al. (2021), and ⃝ Dai et al. (2022) icons represent baselines. The other four icons Refer to caption Refer to caption Refer to caption Refer to caption denote IR2 approaches, indicating the performance of models after fine-tuning on information-regularized synthetic datasets. Metrics are chosen based on standard practice for the three benchmarks. Model performances are averaged across all models used in experiments.
Figure 2: Sample synthetic query from Promptagator and a synthetic query generated with document regularization. (Both queries are generated from the same abstract.) Red indicates overlaps between the Promptagator query and the original abstract. Blue indicates overlaps between the document regularized query and the original abstract. Green indicates overlap with both queries. The document regularized query has less textual overlap with the original query.
This paper focuses on developing methods to generate and utilize synthetic data to enhance the retrieval capabilities of IR systems. To mitigate overfitting, we introduce IR2 (Information Regularization for Information Retrieval), which creates queries that have conceptual overlap with the original document, but varied phrasing and structure. This yields synthetic queries that have a less explicit relationship with the document, see the Document Regularization Query in Figure 2. We posit that training models with such data can be advantageous by nudging them towards understanding deeper semantic relationships, resulting in improved retrieval performance for complex queries.
In this paper, we make several contributions to synthetic data generation for Information Retrieval. First, we introduce three distinct regularization techniques applied at various stages of the synthetic data generation process. This approach addresses models’ tendency to learn superficial and unnatural overlapping features between synthetic queries and documents, which would hamper model performance on complex query IR tasks
We also present a comprehensive analysis of how different pretrained transformer-based embedding models respond to these regularization techniques Wang et al. (2022c); Liu et al. (2019); Cohan et al. (2020). Our empirical evidence demonstrates consistent performance improvements with our methods compared to non-regularized baselines Gao et al. (2021); Dai et al. (2022).
Furthermore, we find that these IR2 regularization techniques can be effectively combined, with instruction and output regularization emerging as the most potent pairing in our tests, as shown by [Uncaptioned image] and [Uncaptioned image] in Figure 1, top right corner. This finding paves the way for more effective strategies in data augmentation for IR systems.
2 Related Work
In recent years, Information Retrieval has seen increasing task complexity as well as advancements in models designed to navigate this greater complexity.
IR tasks have typically centered on retrieving relevant passages to answer simple, sentence-level queries, as seen in foundational datasets like MS MARCO Nguyen et al. (2016) and NQ Kwiatkowski et al. (2019). The BEIR dataset Thakur et al. (2021) combines 19 IR tasks, most of which are sentence-level, including specific challenges like TREC-COVID Voorhees et al. (2021) and SCIFACT Wadden et al. (2020). Several recent datasets contain complex, paragraph-length queries. DORIS-MAE Wang et al. (2023) poses the task of retrieving scientific abstracts given complex scientific research questions. In the WhatsThatBook dataset Lin et al. (2023), tip-of-the-tongue user queries are used to retrieve book descriptions. In ArguAna Wachsmuth et al. (2018), paragraph-length arguments are used to retrieve counterarguments.
Lexicon-based methods such as TF-IDF Sparck Jones (1972) and BM25 Robertson et al. (2009) retrieve based on keyword matching, and are effective in scenarios with substantial token overlap between queries and relevant passages. However, their limitations became apparent as tasks required deeper semantic comprehension, especially in cases with minimal token overlap.
Transformer-based models, including various cross-encoders and dual-encoders Wang et al. (2022c, b); Gao et al. (2021); Cohan et al. (2020); Santhanam et al. (2022); Xiong et al. (2020); Formal et al. (2021), have offered more nuanced document and query representations. They have shown strong results on benchmarks like BEIR Thakur et al. (2021) but have struggled on complex query tasks.
Recently, improvements in LLMs, coupled with prompting techniques such as Chain-of-Thought Kojima et al. (2022); Wei et al. (2022b) and In-Context-Learning Chowdhery et al. (2022); Wei et al. (2022a), have influenced the field of Information Retrieval (IR) in several directions. This has included data annotation Wang et al. (2023); Gilardi et al. (2023); Wang et al. (2022a) as well as retrieval and reranking processes Sun et al. (2023). Most relevant for the current work, LLMs have been used for synthetic data generation, particularly in data-scarce conditions, outperforming traditional augmentation methods Izacard et al. (2022). For instance, Ma et al. (2021); Liang et al. (2020) applied LLMs for synthetic question generation in Question Answering (QA).
In IR tasks, since user-generated queries are frequently difficult or expensive to obtain, much of the work on synthetic data has focused on generating synthetic queries. This has included InPars Bonifacio et al. (2022); Jeronymo et al. (2023) and Promptagator Dai et al. (2022), the latter showcasing significant success on the BEIR benchmark. Augtriever Meng et al. (2022) introduced methods for synthetic query generation using smaller models, optimizing both time and cost. Peng et al. (2023) used soft prompt-tuning to further enhance the quality of generated queries.
Synthetic data generation also has applications in other fields, including text classification Li et al. (2023), information extraction Josifoski et al. (2023), reinforcement learning for language model alignment Yang et al. (2023), human-computer interaction Hämäläinen et al. (2023), and computational social science Veselovsky et al. (2023).
In addition to these LLM developments, there has been an ongoing push to improve sentence embeddings, including SimCSE Gao et al. (2021), which applies dropout masking as a data augmentation technique, and other variants such as DiffCSE Chuang et al. (2022), and RankCSE Liu et al. (2023). By using a contrastive learning objective, these models have substantially increased the utility of unlabeled data for retrieval systems.
3 Methodology
(a) Document Regularization
(b) Instruction Regularization
(c) Query Regularization
Figure 3: Illustration of the three information regularization methods.
3.1 Synthetic Query Generation
The goal of IR systems is to match user queries, which express the user’s information need, with relevant documents that contain the desired information. A typical IR system would benefit from fine-tuning of pairs of query and relevant document if sufficient data is available. However, real user queries are difficult and costly to collect. In addition, finding relevant documents for user queries also requires careful annotations. Therefore, synthetic query generation aims to improve IR system performance when faced with challenges of the scarcity of query data and the lack of supervised pairs of query and relevant document. Since documents are generally more readily available, by generating artificial yet plausible user queries based on existing documents, we can create additional training data for IR systems, helping them better understand and respond to a wider range of potential user queries.
In this research, we concentrate on methods for generating synthetic queries that are relevant to given documents. This involves not just creating queries that a user might realistically pose, but also ensuring that these queries are diverse enough to train IR systems effectively. By using this synthetic data for training, we aim to improve the systems’ overall ability to handle complex user queries, enhancing retrieval accuracy and efficiency.
3.2 Baseline Query Generation
Promptagator Dai et al. (2022) proposes a few-shot method for LLM synthetic query generation. By providing an LLM with a small number of pairs of relevant documents and queries, it will implicitly learn the transformation from document to query. When a new document is provided, the LLM will transform it to a relevant query. We use $d_{i}$ to denote a document and $q_{i}$ to denote a relevant query. Promptagator uses 8-shot prompting of the following form:
$d_{1},q_{1},\dots,d_{8},q_{8},d_{9},$
which includes the 8 example pairs as well as the target document $d_{9}$.
However, as shown in Figure 2, biases from the LLM may result in suboptimal queries. In our testing, we observed that queries frequently contained an implausibly large number of details from the documents, including highly similar phrasing and mirroring structure. When these queries are used for training downstream IR systems, they may lead to overfitting, as the IR systems learn simple phrase-matching heuristics between queries and documents.
3.3 Information Regularization
In order to mitigate these potential problems in the query generation process, we introduce IR2 (Information Regularization for Information Retrieval). These methods aim to reduce the number of superficial features shared by documents and synthetic queries, while still maintaining deeper semantic relationships.
The three information regularization methods can be categorized by the stage of the query generation pipeline at which they are applied:
  • Document Regularization is applied to the input document $d_{i}$ of LLM by intentionally withholding parts of the semantic information in $d_{i}$.
  • Instruction Regularization is applied to the prompt. We design a specialized prompt for the LLM, explicitly instructing it to avoid superficial similarities between the query and document.
  • Query Regularization is applied to the synthetic query which was generated by the LLM. It is designed to summarize the synthetic query, transforming it into a less complex sub-query.
Below we will discuss the motivation and intended effects of each type of regularization.
3.3.1 Input Document Regularization
Document Regularization aims to address implausibly high levels of phrase overlaps between documents and synthetic queries. This technique disrupts phrase overlaps by partially masking the document’s content before feeding it into the LLM. As shown in Figure 3 (a), the system identifies key semantic words in a document and hides a random $p\%$ of them. The LLM, guided by the style of a single example query, then generates a new query based on this incomplete information.
This approach forces the LLM to generate queries that diverge textually from the source document while still remaining broadly relevant, and it makes the resulting training data more challenging for downstream IR systems. The method can be straightforwardly adapted to different IR datasets, and the $p\%$ parameter offers control over the degree of regularization. Setting $p=0$ defaults to the original Promptagator approach, with no regularization.
3.3.2 Instruction Regularization
This method involves guiding the LLM through a structured thought process to generate queries from documents, emphasizing the extraction and synthesis of key ideas while avoiding superficial mimicry. The model is first instructed to break the document into a sequence of aspects, each describing an important part of the document. Each aspect is subsequently paraphrased. The LLM is finally instructed to combine the paraphrased aspects into a natural query. The approach is detailed in Figure 3 (b).
Due to the complexity of the prompt, we found that GPT-3.5 was unable to perform this task, while GPT-4 (which is used for our experiments) was able to.
3.3.3 Output Query Regularization
As depicted in Figure 3 (c), this technique simplifies synthetic queries, making them shorter and less detailed yet still conceptually relevant to the source document. The process challenges embedding models to focus on deeper semantic understanding rather than textual similarity.
Query regularization is straightforward, requires a small context window, and can be combined with our other synthetic generation techniques. In our experiments, we investigate the effectiveness of query regularization applied to synthetic queries generated from document regularization and instruction regularization.
4 Experiments
To evaluate the effectiveness of information regularization for synthetic query generation, we choose 3 complex-query based test benchmarks (DORIS-MAE, ArguAna and WhatsThatBook) and 4 different embedding models (E5-v2-Large Wang et al. (2022c), SimCSE-Large Gao et al. (2021), RoBERTa-Large Liu et al. (2019), SPECTER-v2 Cohan et al. (2020)). For each benchmark, we generate 8 synthetic datasets:
  • Document Regularization ($p=40\%$)
  • Document Regularization ($p=60\%$)
  • Document Regularization ($p=80\%$)
  • Instruction Regularization
  • Query Regularization $\circ$ Document Regularization ($p=40\%$)
  • Query Regularization $\circ$ Document Regularization ($p=60\%$)
  • Query Regularization $\circ$ Document Regularization ($p=80\%$)
  • Query Regularization $\circ$ Instruction Regularization.
The notation $\circ$ indicates function composition.
In this section, we will discuss the details of benchmark datasets, pretrained models, baseline query generation methods, and regularization techniques. All experiment code is publicly accessible. [^1] https://github.com/Info-Regularization/Information-Regularization
For model evaluations, we report the IR metrics recommended by each benchmark’s associated paper. Full experimental results with a more complete set of metrics are reported in Appendix B.
4.1 Benchmarks
WhatsThatBook (WTB) Lin et al. (2023) contains 14,441 queries from users trying to recall specific books, paired with book titles and descriptions. The task is to match each of the 1,445 test set queries with the correct descriptions. We use descriptions from 4,000 unique books, exclusive from the test set, for synthetic query generation.
DORIS-MAE Wang et al. (2023) includes 100 complex research queries in AI, CV, NLP, and ML, split into 40 for training and 60 for testing. Each query is associated with a candidate pool of approximately 100 research paper abstracts, with a fine-grained ranking system. The candidate pools are drawn from a corpus of approximately 360,000 computer science papers. We use 4,000 abstracts from a 360,000 paper corpus for synthetic query generation.
ArguAna Wachsmuth et al. (2018) from BEIR Thakur et al. (2021) consists of 8,674 paragraph-length arguments, and a test set of 1,406 argument-counterargument pairs. The task is to retrieve the correct counterargument for each test set argument. We use 5,700 arguments, not part of the test set, for synthetic counterargument generation. Due to the symmetry between arguments and counterarguments, we can use our synthetic data to train for retrieval of counterarguments given arguments.
4.2 Models
We experiment with four models: E5-Large-v2 Wang et al. (2022c) (355M), SimCSE-Large Gao et al. (2021) (355M), RoBERTa-Large Liu et al. (2019) (355M), and SPECTER-v2 Cohan et al. (2020) (110M). E5-Large-v2 (denoted as E5), a contrastively trained model, demonstrates performance on par with the state-of-the-art ada-002 Greene et al. (2022) on BEIR Thakur et al. (2021). For SimCSE-Large (denoted as SimCSE), which uses self-supervised training with dropout masks, we choose the top-performing checkpoint on the STS benchmark Agirre et al. (2013). RoBERTa-Large is denoted as RoBERTa. SPECTER-v2 is specialized for scientific document representation. Embeddings are derived according to each model’s specifications, using pooling strategies, and retrieval and reranking are based on similarity measures (L2 distance for SPECTER-v2; cosine similarity for others).
4.3 Baselines
Three baseline types were used. First, the "Pretrained" baselines indicate performance of the pretrained checkpoints on each dataset. Second, a "Contrastive Fine-tuning" approach, similar to SimCSE Gao et al. (2021), applies random dropout masks to documents, creating self-supervised pairs for fine-tuning all models. This is used as a domain adaptation baseline. Third, we compare to Promptagator Dai et al. (2022), a few-shot synthetic query generation method. Its synthetic data generation process was replicated, using GPT-4 (gpt-4-0613) instead of FLAN 137B Wei et al. (2021).
To eliminate possible data confounds, all baselines use the same data as the regularization methods. For each benchmark, the regularization methods use a subset of documents for synthetic query generation. For the Promptagator baseline, the same documents are used for synthetic generation. For the Contrastive Fine-tuning baseline, the embedding models are directly fine-tuned on the same documents. Specifically, we train using 4000 documents from DORIS-MAE’s corpus, 5700 arguments from ArguAna, and 4000 book descriptions from WhatsThatBook. We removed documents which were unusually short or long, and randomly sampled from the remaining documents.
E5-Large-v2 RoBERTa
Method R@5 R@20 NDCG@10 MRR@10 R@5 R@20 NDCG@10 MRR@10
Pretrained 18.41 27.06 15.25 13.07 2.84 5.33 2.17 1.73
Contrastive Fine-tuning 37.25 49.85 33.16 29.87 23.59 35.27 20.69 18.10
Promptagator 41.53 54.66 37.07 33.56 27.23 39.14 23.69 20.78
Doc40% reg 41.40 53.69 $\downarrow$ -0.97 37.01 33.72 28.30 $\uparrow$ 1.07 40.89 $\uparrow$ 1.75 24.79 $\uparrow$ 1.10 21.81 $\uparrow$ 1.03
Queryreg $\circ$ Doc40% reg 45.50 $\uparrow$ 3.97 58.08 $\uparrow$ 3.42 40.50 $\uparrow$ 3.43 37.01 $\uparrow$ 3.45 32.83 $\uparrow$ 5.60 46.38 $\uparrow$ 7.24 29.05 $\uparrow$ 5.36 25.82 $\uparrow$ 5.04
Doc60% reg 44.15 $\uparrow$ 2.62 55.69 $\uparrow$ 1.03 39.64 $\uparrow$ 2.57 36.16 $\uparrow$ 2.60 28.67 $\uparrow$ 1.44 41.49 $\uparrow$ 2.35 25.14 $\uparrow$ 1.45 22.06 $\uparrow$ 1.28
Queryreg $\circ$ Doc60% reg 46.23 $\uparrow$ 4.70 59.76 $\uparrow$ 5.10 41.58 $\uparrow$ 4.51 38.13 $\uparrow$ 4.57 34.30 $\uparrow$ 7.07 47.52 $\uparrow$ 8.38 30.15 $\uparrow$ 6.46 26.83 $\uparrow$ 6.05
Doc80% reg 28.30 $\downarrow$ -13.23 39.85 $\downarrow$ -14.81 24.05 $\downarrow$ -13.02 21.13 $\downarrow$ -12.43 21.74 $\downarrow$ -5.49 33.16 $\downarrow$ -5.98 18.91 $\downarrow$ -4.78 16.35 $\downarrow$ -4.43
Queryreg $\circ$ Doc80% reg 44.80 $\uparrow$ 3.27 58.21 $\uparrow$ 3.55 40.29 $\uparrow$ 3.22 36.94 $\uparrow$ 3.38 33.69 $\uparrow$ 6.46 46.23 $\uparrow$ 7.09 29.41 $\uparrow$ 5.72 26.20 $\uparrow$ 5.42
Instrreg 42.15 $\uparrow$ 0.62 53.02 $\downarrow$ -1.64 37.33 33.89 $\uparrow$ 0.33 28.12 $\uparrow$ 0.89 40.80 $\uparrow$ 1.66 24.51 $\uparrow$ 0.82 21.46 $\uparrow$ 0.68
Queryreg $\circ$ Instrreg 44.90 $\uparrow$ 3.37 56.28 $\uparrow$ 1.62 39.29 $\uparrow$ 2.22 35.79 $\uparrow$ 2.23 33.09 $\uparrow$ 5.86 46.09 $\uparrow$ 6.95 29.28 $\uparrow$ 5.59 26.06 $\uparrow$ 5.28
SimCSE SPECTER-v2
Method R@5 R@20 NDCG@10 MRR@10 R@5 R@20 NDCG@10 MRR@10
Pretrained 16.96 24.98 14.58 12.67 2.98 5.26 2.43 1.87
Contrastive Fine-tuning 26.78 37.15 23.11 20.42 7.85 12.56 6.36 5.35
Promptagator 30.97 44.26 27.15 23.99 6.56 11.81 5.72 4.88
Doc40% reg 32.10 $\uparrow$ 1.13 43.07 $\downarrow$ -1.19 28.30 $\uparrow$ 1.15 25.45 $\uparrow$ 1.46 9.98 $\uparrow$ 3.42 16.12 $\uparrow$ 4.31 8.55 $\uparrow$ 2.83 7.32 $\uparrow$ 2.44
Queryreg $\circ$ Doc40% reg 34.63 $\uparrow$ 3.66 47.44 $\uparrow$ 3.18 30.64 $\uparrow$ 3.49 27.51 $\uparrow$ 3.52 10.84 $\uparrow$ 4.28 16.26 $\uparrow$ 4.45 8.81 $\uparrow$ 3.09 7.57 $\uparrow$ 2.69
Doc60% reg 33.03 $\uparrow$ 2.06 43.96 28.87 $\uparrow$ 1.72 25.97 $\uparrow$ 1.98 8.64 $\uparrow$ 2.08 14.49 $\uparrow$ 2.68 7.46 $\uparrow$ 1.74 6.26 $\uparrow$ 1.38
Queryreg $\circ$ Doc60% reg 34.84 $\uparrow$ 3.87 47.08 $\uparrow$ 2.82 30.77 $\uparrow$ 3.62 27.64 $\uparrow$ 3.65 7.54 $\uparrow$ 0.98 13.29 $\uparrow$ 1.48 6.55 $\uparrow$ 0.83 5.48 $\uparrow$ 0.60
Doc80% reg 19.75 $\downarrow$ -11.22 29.89 $\downarrow$ -14.37 16.73 $\downarrow$ -10.42 14.33 $\downarrow$ -9.66 7.39 $\uparrow$ 0.83 13.19 $\uparrow$ 1.38 6.46 $\uparrow$ 0.74 5.31 $\uparrow$ 0.43
Queryreg $\circ$ Doc80% reg 33.76 $\uparrow$ 2.79 45.63 $\uparrow$ 1.37 29.34 $\uparrow$ 2.19 26.26 $\uparrow$ 2.27 8.14 $\uparrow$ 1.58 14.69 $\uparrow$ 2.88 7.33 $\uparrow$ 1.61 6.13 $\uparrow$ 1.25
Instrreg 33.81 $\uparrow$ 2.84 45.79 $\uparrow$ 1.53 29.74 $\uparrow$ 2.59 26.79 $\uparrow$ 2.80 10.49 $\uparrow$ 3.93 16.80 $\uparrow$ 4.99 9.30 $\uparrow$ 3.58 8.07 $\uparrow$ 3.19
Queryreg $\circ$ Instrreg 37.34 $\uparrow$ 6.37 50.45 $\uparrow$ 6.19 33.18 $\uparrow$ 6.03 30.02 $\uparrow$ 6.03 11.08 $\uparrow$ 4.52 17.36 $\uparrow$ 5.55 9.59 $\uparrow$ 3.87 8.28 $\uparrow$ 3.40
Table 1: Results for WhatsThatBook. The average of 20 random trials is reported for each model/method/metric. Green arrow indicates a statistically significant ($p<0.05$) increase over Promptagator baseline. Red arrow indicates a significant decrease.
4.4 Implementation of Regularization
Input Document Regularization: We use gpt-3.5-turbo-0301 to extract keywords/phrases from a target document, subsequently redacting $p\%$ of them. gpt-4-0613 then generates a query from this redacted document, guided solely by an example query format, without paired documents. (For ArguAna, we do not provide an example query format, since the LLM has sufficient prior knowledge of the argument and counterargument format.)
Prompt Instruction Regularization: Refer to Figure 3 (b) and Section 3.3.2. We use gpt-4-0613 with a prompt designed to elicit the breakdown and paraphrasing of query aspects. All prompts are available in the project GitHub repository. [^2] https://github.com/Info-Regularization/Information-Regularization/tree/main/prompt
Output Query Regularization: gpt-3.5-turbo-0301 is used to summarize the original query into a simpler query. This method was applied to both document and instruction regularization outputs in our tests.
For all prompts used for keyword extraction and synthetic query generation, see Appendix A.
4.5 Training Methods
We fine-tuned each model on the different regularized datasets for one epoch, utilizing the AdamW optimizer Loshchilov and Hutter (2017) with an initial learning rate of 1e-5, coupled with a cosine annealing schedule. While various warm-up ratios were tested, omitting the warm-up phase yielded more stable training/validation loss curves and enhanced performance. No additional hyperparameter tuning was done; hyperparameters were set to their default values. To maintain consistent training memory, texts exceeding 512 tokens were truncated to their initial 512 tokens, a sufficient range as most texts fell within this limit. We employed the standard NT-Xent loss with in-batch negatives, setting the temperature $\tau=0.05$ and a batch size of 80, the maximum capacity for four 40G NVIDIA A100 GPUs. To minimize gradient noise during minibatch processing, we accumulated gradients across 20 minibatches.
The synthetically generated query $q_{i}$, the relevant document $d_{i}$, and the batch size $N$ are used as inputs to an embedding function $f$ with range in $\mathbb{R}^{n}$. The NT-Xent Loss Chen et al. (2020) is then defined as:
$L=\underset{{1\leq i\leq N}}{\mathbb{E}}\left[-\log\frac{e^{\langle f(q_{i}),f(d_{i})\rangle/\tau}}{\underset{1\leq j\leq N}{\sum}e^{\langle f(q_{i}),f(d_{j})\rangle/\tau}}\right]$ (1)
All experiments were performed using 20 random seeds, with the averaged results shown in Tables 1, 2 and 3. We use t-tests to compare the effectiveness of our regularization strategies against the 8-shot Promptagator baseline. Because multiple methods are being compared, we use Bonferroni correction Weisstein (2004) with a correction factor 8 (the number of IR2 methods). Using the adjusted p-value, colored arrows in the tables indicate statistically significant increases or decreases (adjusted $p<0.05$) relative to the Promptagator baseline.
E5-Large-v2 RoBERTa
Method R@5 R@20 NDCG@10 MRR@10 R@5 R@20 NDCG@10 MRR@10
Pretrained 14.67 42.15 71.98 14.34 11.96 34.77 66.86 8.56
Contrastive Fine-tuning 15.28 39.87 72.51 13.43 14.75 43.32 73.55 20.79
Promptagator 16.18 45.59 73.95 14.83 13.03 42.88 72.23 16.82
Doc40% reg 14.90 $\downarrow$ -1.28 46.55 $\uparrow$ 0.96 74.28 $\uparrow$ 0.33 14.87 12.54 44.89 $\uparrow$ 2.01 72.94 $\uparrow$ 0.71 18.16 $\uparrow$ 1.34
Queryreg $\circ$ Doc40% reg 15.85 45.67 73.96 16.10 $\uparrow$ 1.27 14.66 $\uparrow$ 1.63 47.41 $\uparrow$ 4.53 74.04 $\uparrow$ 1.81 18.81 $\uparrow$ 1.99
Doc60% reg 15.23 $\downarrow$ -0.95 46.57 $\uparrow$ 0.98 74.98 $\uparrow$ 1.03 15.14 14.11 $\uparrow$ 1.08 43.99 $\uparrow$ 1.11 72.94 $\uparrow$ 0.71 16.84
Queryreg $\circ$ Doc60% reg 15.31 $\downarrow$ -0.87 48.16 $\uparrow$ 2.57 74.90 $\uparrow$ 0.95 15.43 $\uparrow$ 0.60 14.75 $\uparrow$ 1.72 47.25 $\uparrow$ 4.37 74.61 $\uparrow$ 2.38 20.48 $\uparrow$ 3.66
Doc80% reg 14.68 $\downarrow$ -1.50 43.89 $\downarrow$ -1.70 74.13 $\uparrow$ 0.18 13.51 $\downarrow$ -1.32 14.47 $\uparrow$ 1.44 42.47 72.61 $\uparrow$ 0.38 15.10 $\downarrow$ -1.72
Queryreg $\circ$ Doc80% reg 15.63 $\downarrow$ -0.55 46.75 $\uparrow$ 1.16 75.08 $\uparrow$ 1.13 16.62 $\uparrow$ 1.79 15.84 $\uparrow$ 2.81 46.96 $\uparrow$ 4.08 74.71 $\uparrow$ 2.48 20.49 $\uparrow$ 3.67
Instrreg 16.38 46.73 $\uparrow$ 1.14 74.42 $\uparrow$ 0.47 17.06 $\uparrow$ 2.23 13.80 $\uparrow$ 0.77 45.78 $\uparrow$ 2.90 73.53 $\uparrow$ 1.30 18.24 $\uparrow$ 1.42
Queryreg $\circ$ Instrreg 15.76 $\downarrow$ -0.42 48.46 $\uparrow$ 2.87 76.02 $\uparrow$ 2.07 20.86 $\uparrow$ 6.03 15.20 $\uparrow$ 2.17 47.44 $\uparrow$ 4.56 74.08 $\uparrow$ 1.85 19.31 $\uparrow$ 2.49
SimCSE SPECTER-v2
Method R@5 R@20 NDCG@10 MRR@10 R@5 R@20 NDCG@10 MRR@10
Pretrained 14.38 41.83 70.81 24.83 13.59 41.92 71.46 21.74
Contrastive Fine-tuning 15.60 44.99 73.18 22.41 14.56 45.64 71.46 20.49
Promptagator 16.35 46.69 73.97 16.54 14.13 39.22 71.55 22.37
Doc40% reg 17.70 $\uparrow$ 1.35 45.30 $\downarrow$ -1.39 74.49 $\uparrow$ 0.52 16.49 15.39 $\uparrow$ 1.26 43.39 $\uparrow$ 4.17 72.11 $\uparrow$ 0.56 21.66 $\downarrow$ -0.71
Queryreg $\circ$ Doc40% reg 16.06 45.46 $\downarrow$ -1.23 73.23 $\downarrow$ -0.74 16.56 14.19 42.66 $\uparrow$ 3.44 71.49 21.40 $\downarrow$ -0.97
Doc60% reg 17.08 $\uparrow$ 0.73 45.33 $\downarrow$ -1.36 74.70 $\uparrow$ 0.73 17.75 $\uparrow$ 1.21 15.41 $\uparrow$ 1.28 44.19 $\uparrow$ 4.97 72.81 $\uparrow$ 1.26 21.35 $\downarrow$ -1.02
Queryreg $\circ$ Doc60% reg 17.04 $\uparrow$ 0.69 46.68 74.65 $\uparrow$ 0.68 18.32 $\uparrow$ 1.78 14.95 $\uparrow$ 0.82 43.30 $\uparrow$ 4.08 72.19 $\uparrow$ 0.64 21.60 $\downarrow$ -0.77
Doc80% reg 15.01 $\downarrow$ -1.34 42.53 $\downarrow$ -4.16 73.69 $\downarrow$ -0.28 18.68 $\uparrow$ 2.14 17.02 $\uparrow$ 2.89 44.97 $\uparrow$ 5.75 73.94 $\uparrow$ 2.39 22.78 $\uparrow$ 0.41
Queryreg $\circ$ Doc80% reg 15.92 45.70 $\downarrow$ -0.99 73.86 18.54 $\uparrow$ 2.00 14.70 $\uparrow$ 0.57 44.66 $\uparrow$ 5.44 73.07 $\uparrow$ 1.52 22.65
Instrreg 16.50 46.68 75.60 $\uparrow$ 1.63 19.71 $\uparrow$ 3.17 14.73 $\uparrow$ 0.60 41.50 $\uparrow$ 2.28 73.06 $\uparrow$ 1.51 21.05 $\downarrow$ -1.32
Queryreg $\circ$ Instrreg 16.06 47.61 $\uparrow$ 0.92 75.43 $\uparrow$ 1.46 22.14 $\uparrow$ 5.60 14.50 $\uparrow$ 0.37 41.44 $\uparrow$ 2.22 72.46 $\uparrow$ 0.91 22.26
Table 2: Result for DORIS-MAE. See Table 1 caption for reporting conventions.
5 Results
5.1 Results on WhatsThatBook
Table 1 shows comparisons between the IR2 methods and the Promptagator baseline. Except for Doc80% reg, RoBERTa and SPECTER-v2 show consistent improvement with all regularization methods. For E5 and SimCSE, there is no clear signal of improvement from document or instruction regularization alone. However, query regularization, combined with either document or instruction regularization, improves performance across all models.
Not all models respond to fine-tuning on synthetic datasets in the same way. SPECTER-v2 and RoBERTa start from similar performance levels, but RoBERTa reaches considerably higher performance after fine-tuning. This suggests that the effectiveness of synthetic data augmentation is model-dependent. The SPECTER-v2 checkpoint had been previously fine-tuned for scientific document understanding, possibly interfering with its performance on the WhatsThatBook dataset.
Document regularization with an 80% masking ratio leads to synthetic queries that are either too vague or contain misleading and false information. Consequently, for E5, RoBERTa, and SimCSE, we observed that document regularization with an 80% mask resulted in significantly lower performances compared to Promptagator. Results are stronger with a masking level between 40% and 60%.
5.2 Results on DORIS-MAE
Table 2 shows that instruction regularization, alone or combined with query regularization, consistently enhances performance across various models and metrics, surpassing the Promptagator’s synthetic dataset, with minor exceptions within a 1% margin. The strongest results across all models and data synthesis methods are achieved by query regularization combined with instruction regularization.
Document regularization has variable performance depending on the mask percentage ($p=40\%,60\%,80\%$) and model. It has only a small effect on the performance of E5-Large-v2 compared to the Promptagator baseline, and underperforms with SimCSE. However, it substantially improves performance when applied to RoBERTa and SPECTER-v2.
For E5, SimCSE, and RoBERTa, integrating query regularization with document or instruction regularization improves performance across all metrics. Query regularization decreases the textual similarity between queries and their target documents, suggesting that models are learning deeper semantic relationships between queries and documents.
5.3 Results on ArguAna
E5-Large-v2 RoBERTa
Method R@5 R@20 NDCG@10 MRR@10 R@5 R@20 NDCG@10 MRR@10
Pretrained 59.37 88.16 47.75 38.96 30.51 49.68 23.56 18.49
Contrastive Fine-tuning 70.69 94.77 56.15 47.13 59.12 84.67 47.57 39.65
Promptagator 69.41 94.11 55.14 46.67 61.51 88.19 49.80 41.85
Doc40% reg 78.72 $\uparrow$ 9.31 96.46 $\uparrow$ 2.35 64.24 $\uparrow$ 9.10 56.45 $\uparrow$ 9.78 67.38 $\uparrow$ 5.87 91.03 $\uparrow$ 2.84 54.54 $\uparrow$ 4.74 46.43 $\uparrow$ 4.58
Queryreg $\circ$ Doc40% reg 77.76 $\uparrow$ 8.35 96.21 $\uparrow$ 2.10 63.00 $\uparrow$ 7.86 55.24 $\uparrow$ 8.57 67.29 $\uparrow$ 5.78 91.08 $\uparrow$ 2.89 54.04 $\uparrow$ 4.24 45.85 $\uparrow$ 4.00
Doc60% reg 78.16 $\uparrow$ 8.75 96.41 $\uparrow$ 2.30 64.41 $\uparrow$ 9.27 56.57 $\uparrow$ 9.90 67.51 $\uparrow$ 6.00 90.77 $\uparrow$ 2.58 54.85 $\uparrow$ 5.05 46.85 $\uparrow$ 5.00
Queryreg $\circ$ Doc60% reg 78.34 $\uparrow$ 8.93 96.37 $\uparrow$ 2.26 64.10 $\uparrow$ 8.96 56.24 $\uparrow$ 9.57 67.71 $\uparrow$ 6.20 90.99 $\uparrow$ 2.80 54.41 $\uparrow$ 4.61 46.34 $\uparrow$ 4.49
Doc80% reg 77.37 $\uparrow$ 7.96 95.57 $\uparrow$ 1.46 63.14 $\uparrow$ 8.00 55.16 $\uparrow$ 8.49 67.82 $\uparrow$ 6.31 90.18 $\uparrow$ 1.99 54.62 $\uparrow$ 4.82 46.79 $\uparrow$ 4.94
Queryreg $\circ$ Doc80% reg 78.44 $\uparrow$ 9.03 96.25 $\uparrow$ 2.14 64.36 $\uparrow$ 9.22 56.49 $\uparrow$ 9.82 67.38 $\uparrow$ 5.87 90.70 $\uparrow$ 2.51 54.53 $\uparrow$ 4.73 46.59 $\uparrow$ 4.74
Instrreg 72.06 $\uparrow$ 2.65 95.70 $\uparrow$ 1.59 57.76 $\uparrow$ 2.62 49.17 $\uparrow$ 2.50 64.47 $\uparrow$ 2.96 90.09 $\uparrow$ 1.90 51.49 $\uparrow$ 1.69 43.20 $\uparrow$ 1.35
Queryreg $\circ$ Instrreg 71.59 $\uparrow$ 2.18 95.18 $\uparrow$ 1.07 56.89 $\uparrow$ 1.75 48.37 $\uparrow$ 1.70 65.14 $\uparrow$ 3.63 90.26 $\uparrow$ 2.07 52.25 $\uparrow$ 2.45 43.89 $\uparrow$ 2.04
SimCSE SPECTER-v2
Method R@5 R@20 NDCG@10 MRR@10 R@5 R@20 NDCG@10 MRR@10
Pretrained 49.03 81.12 39.23 30.35 38.91 71.79 30.91 23.23
Contrastive Fine-tuning 36.70 64.79 28.85 22.12 41.90 75.70 33.20 25.26
Promptagator 69.49 92.54 56.38 47.97 38.64 75.15 30.24 21.77
Doc40% reg 74.92 $\uparrow$ 5.43 93.36 $\uparrow$ 0.82 60.24 $\uparrow$ 3.86 52.42 $\uparrow$ 4.45 56.09 $\uparrow$ 17.45 86.32 $\uparrow$ 11.17 45.00 $\uparrow$ 14.76 36.30 $\uparrow$ 14.53
Queryreg $\circ$ Doc40% reg 76.07 $\uparrow$ 6.58 93.97 $\uparrow$ 1.43 61.38 $\uparrow$ 5.00 53.57 $\uparrow$ 5.60 51.28 $\uparrow$ 12.64 83.33 $\uparrow$ 8.18 40.59 $\uparrow$ 10.35 31.81 $\uparrow$ 10.04
Doc60% reg 73.20 $\uparrow$ 3.71 92.10 58.46 $\uparrow$ 2.08 50.58 $\uparrow$ 2.61 58.87 $\uparrow$ 20.23 86.93 $\uparrow$ 11.78 46.55 $\uparrow$ 16.31 37.88 $\uparrow$ 16.11
Queryreg $\circ$ Doc60% reg 75.75 $\uparrow$ 6.26 93.19 $\uparrow$ 0.65 60.94 $\uparrow$ 4.56 53.24 $\uparrow$ 5.27 54.22 $\uparrow$ 15.58 84.97 $\uparrow$ 9.82 42.90 $\uparrow$ 12.66 34.21 $\uparrow$ 12.44
Doc80% reg 70.38 89.67 $\downarrow$ -2.87 55.76 47.97 61.02 $\uparrow$ 22.38 87.48 $\uparrow$ 12.33 48.68 $\uparrow$ 18.44 40.24 $\uparrow$ 18.47
Queryreg $\circ$ Doc80% reg 74.10 $\uparrow$ 4.61 92.59 59.79 $\uparrow$ 3.41 52.18 $\uparrow$ 4.21 57.55 $\uparrow$ 18.91 86.65 $\uparrow$ 11.50 46.01 $\uparrow$ 15.77 37.44 $\uparrow$ 15.67
Instrreg 70.97 93.91 $\uparrow$ 1.37 57.41 48.80 43.26 $\uparrow$ 4.62 80.16 $\uparrow$ 5.01 34.80 $\uparrow$ 4.56 26.07 $\uparrow$ 4.30
Queryreg $\circ$ Instrreg 71.46 $\uparrow$ 1.97 93.78 $\uparrow$ 1.24 58.16 $\uparrow$ 1.78 49.78 $\uparrow$ 1.81 38.57 77.12 $\uparrow$ 1.97 31.72 $\uparrow$ 1.48 23.15 $\uparrow$ 1.38
Table 3: Results for ArguAna. See Table 1 caption for reporting conventions.
In Table 3, we observe that the regularized synthetic datasets consistently outperform baseline approaches. While instruction regularization continues to outperform baselines, all models achieve the strongest performance when fine-tuning on document regularized dataset. Similar to our observations in DORIS-MAE and WTB, query regularization, when combined with another synthetic query method, remains consistently strong.
5.4 Analysis and Discussion
Our experiments demonstrate substantial improvements from regularizing synthetic queries in IR tasks. Rather than simply altering task difficulty, these methods refine the model’s ability to discern and process relevant information, an advancement over existing strategies. Specifically, integrating query and instruction regularization emerges as a robust method, enhancing the model’s performance consistently.
Models trained with the proposed regularized synthetic data generation strategies not only outperform pretrained counterparts but also exhibit considerable gains over existing synthetic data methods like Promptagator. This is shown across various metrics and models.
The practical implications of these methods extend beyond improved model performance. They point toward a greater adaptability in handling complex information retrieval tasks, a critical component given the variability of real-world applications. Moreover, the insights gained from this research contribute to the broader discourse on the role of synthetic data in model training, especially in scenarios where data scarcity is a challenge.
Though IR2 improves over prior methods, there are several caveats. The document regularization method, which masks part of the input document and generates a query from this masked document, sometimes results in hallucinations in the queries. We hypothesize that performance differences across datasets are due to varying task sensitivity to these hallucinations. Performance was generally weaker with a mask ratio above 60%.
As shown in Table 1, document regularization had weaker performance on the WhatsThatBook dataset. This task requires a match between fine-grained details in the query and document. In contrast, for the ArguAna dataset in Table 3, where the task involves pairing arguments with counterarguments, the method had stronger performance. Thematic similarity is sufficient for this task.
5.5 Cost Analysis
Instruction Regularization, which costs \$100-200 per dataset, is the most expensive of the three regularization methods. Due to longer prompts with the inclusion of 8 example pairs, the baseline method Promptagator costs \$400-\$600 per dataset.
We attempted to generate synthetic queries by hand to estimate human labor costs, and required at least 5 minutes per query. Assuming a wage of \$15 per hour, human generation costs at least \$5,000 per dataset. This suggests substantial savings by using GPT-4.
6 Limitations
The effectiveness of our regularization strategies does show variability across different models, signaling that these methods may be more suited to certain architectures or training setups. Future work could delve deeper into this aspect, aiming to identify the underpinning factors that contribute to these discrepancies.
Appendix A Prompt Design
A.1 DORIS-MAE Prompt
Note, Query Regularization for DORIS-MAE will generate 4 shorter synthetic queries per document, we randomly choose one for later training.
  • Figure 4 for Instruction Regularization prompt
  • Figure 5 for Query Regularization on Instruction Regularization prompt
  • Figure 6 for Document Regularization 40%, 60% prompt
  • Figure 7 for Document Regularization 80% prompt
  • Figure 8 for Query Regularization on synthetic queries from Doc Reg 40%, 60%, and 80%
  • Figure 9 for keywords extraction prompt
  • Figure 10 for Promptagator prompt example
Figure 4: DORIS-MAE Instruction Regularization Prompt
Figure 5: DORIS-MAE Query Regularization on Instruction Regularization Prompt
Figure 6: DORIS-MAE Document Regularization 40%, 60% Prompt
Figure 7: DORIS-MAE Document Regularization 80% Prompt
Figure 8: DORIS-MAE Query Regularization on synthetic queries from Doc Reg 40%, 60%, and 80%
Figure 9: DORIS-MAE Keywords Extraction Prompt
Figure 10: DORIS-MAE Promptagator Prompt
A.2 ArguAna Prompt
  • Figure 11 for Instruction Regularization prompt
  • Figure 12 for Query Regularization on Instruction Regularization prompt
  • Figure 13 for Document Regularization 40%, 60%, and 80% prompt
  • Figure 14 for Query Regularization on synthetic queries from Doc Reg 40%, 60%, and 80%
  • Figure 15 for keywords extraction prompt
  • Figure 16 for Promptagator prompt example
Figure 11: ArguAna Instruction Regularization Prompt
Figure 12: ArguAna Query Regularization on Instruction Regularization Prompt
Figure 13: ArguAna Document Regularization 40%, 60%, and 80% Prompt
Figure 14: ArguAna Query Regularization on synthetic queries from Doc Reg 40%, 60%, and 80%
Figure 15: ArguAna Keywords Extraction Prompt
Figure 16: ArguAna Promptagator Prompt
A.3 WhatsThatBook Prompt
  • Figure 17 for Instruction Regularization prompt
  • Figure 18 for Query Regularization on Instruction Regularization prompt
  • Figure 19 for Document Regularization 40% prompt
  • Figure 20 for Document Regularization 60%, 80% prompt
  • Figure 21 for Query Regularization on synthetic queries from Doc Reg 40%, 60%, and 80%
  • Figure 22 for keywords extraction prompt
  • Figure 23 for Promptagator prompt example
Figure 17: WhatsThatBook Instruction Regularization Prompt
Figure 18: WhatsThatBook Query Regularization on Instruction Regularization Prompt
Figure 19: WhatsThatBook Document Regularization 40% Prompt
Figure 20: WhatsThatBook Document regularization 60%, 80% Prompt
Figure 21: WhatsThatBook Query Regularization on synthetic queries from Doc Reg 40%, 60%, and 80%
Figure 22: WhatsThatBook Keywords Extraction Prompt
Figure 23: WhatsThatBook Promptagator Prompt
Appendix B Experiment Details
We provide the experimental results on the three datasets in Table 4, 5, 6 with a more complete set of metrics. The results are consistent with those reported in Section 5.
Model (Method) R@5 R@10 R@20 RP NDCG@10 MRR@10 MAP
E5-Large-v2
Pretrained 14.67 25.98 42.15 38.18 71.98 14.34 40.52
Contrastive Fine-tuning 15.28 25.97 39.87 34.73 72.51 13.43 38.87
Promptagator 16.18 26.58 45.59 38.99 73.95 14.83 42.69
Document40% reg 14.90 $\downarrow$ -1.28 27.65 $\uparrow$ 1.07 46.55 $\uparrow$ 0.96 39.33 74.28 $\uparrow$ 0.33 14.87 43.10 $\uparrow$ 0.41
Queryreg $\circ$ Document40% reg 15.85 27.44 $\uparrow$ 0.86 45.67 39.73 $\uparrow$ 0.74 73.96 16.10 $\uparrow$ 1.27 43.06 $\uparrow$ 0.37
Document60% reg 15.23 $\downarrow$ -0.95 27.68 $\uparrow$ 1.10 46.57 $\uparrow$ 0.98 38.86 74.98 $\uparrow$ 1.03 15.14 43.26 $\uparrow$ 0.57
Queryreg $\circ$ Document60% reg 15.31 $\downarrow$ -0.87 27.74 $\uparrow$ 1.16 48.16 $\uparrow$ 2.57 39.60 $\uparrow$ 0.61 74.90 $\uparrow$ 0.95 15.43 $\uparrow$ 0.60 44.48 $\uparrow$ 1.79
Document80% reg 14.68 $\downarrow$ -1.50 25.02 $\downarrow$ -1.56 43.89 $\downarrow$ -1.70 38.91 74.13 $\uparrow$ 0.18 13.51 $\downarrow$ -1.32 43.31 $\uparrow$ 0.62
Queryreg $\circ$ Document80% reg 15.63 $\downarrow$ -0.55 27.93 $\uparrow$ 1.35 46.75 $\uparrow$ 1.16 39.37 75.08 $\uparrow$ 1.13 16.62 $\uparrow$ 1.79 44.24 $\uparrow$ 1.55
Instrreg 16.38 26.26 46.73 $\uparrow$ 1.14 40.31 $\uparrow$ 1.32 74.42 $\uparrow$ 0.47 17.06 $\uparrow$ 2.23 44.84 $\uparrow$ 2.15
Queryreg $\circ$ Instrreg 15.76 $\downarrow$ -0.42 30.56 $\uparrow$ 3.98 48.46 $\uparrow$ 2.87 42.25 $\uparrow$ 3.26 76.02 $\uparrow$ 2.07 20.86 $\uparrow$ 6.03 46.29 $\uparrow$ 3.60
SimCSE
Pretrained 14.38 23.24 41.83 36.27 70.81 24.83 42.02
Contrastive Fine-tuning 15.60 27.70 44.99 39.06 73.18 22.41 42.87
Promptagator 16.35 28.98 46.69 40.64 73.97 16.54 44.12
Document40% reg 17.70 $\uparrow$ 1.35 29.85 $\uparrow$ 0.87 45.30 $\downarrow$ -1.39 40.43 74.49 $\uparrow$ 0.52 16.49 44.67 $\uparrow$ 0.55
Queryreg $\circ$ Document40% reg 16.06 28.77 45.46 $\downarrow$ -1.23 40.41 73.23 $\downarrow$ -0.74 16.56 43.70 $\downarrow$ -0.42
Document60% reg 17.08 $\uparrow$ 0.73 28.80 45.33 $\downarrow$ -1.36 39.91 $\downarrow$ -0.73 74.70 $\uparrow$ 0.73 17.75 $\uparrow$ 1.21 44.50 $\uparrow$ 0.38
Queryreg $\circ$ Document60% reg 17.04 $\uparrow$ 0.69 28.63 46.68 40.59 74.65 $\uparrow$ 0.68 18.32 $\uparrow$ 1.78 45.05 $\uparrow$ 0.93
Document80% reg 15.01 $\downarrow$ -1.34 26.37 $\downarrow$ -2.61 42.53 $\downarrow$ -4.16 38.32 $\downarrow$ -2.32 73.69 $\downarrow$ -0.28 18.68 $\uparrow$ 2.14 42.22 $\downarrow$ -1.90
Queryreg $\circ$ Document80% reg 15.92 27.32 $\downarrow$ -1.66 45.70 $\downarrow$ -0.99 39.59 $\downarrow$ -1.05 73.86 18.54 $\uparrow$ 2.00 43.96
Instrreg 16.50 27.81 $\downarrow$ -1.17 46.68 40.71 75.60 $\uparrow$ 1.63 19.71 $\uparrow$ 3.17 45.85 $\uparrow$ 1.73
Queryreg $\circ$ Instrreg 16.06 28.25 $\downarrow$ -0.73 47.61 $\uparrow$ 0.92 39.88 $\downarrow$ -0.76 75.43 $\uparrow$ 1.46 22.14 $\uparrow$ 5.60 45.52 $\uparrow$ 1.40
RoBERTa
Pretrained 11.96 21.45 34.77 30.68 66.86 8.56 34.14
Contrastive Fine-tuning 14.75 25.95 43.32 37.68 73.55 20.79 41.67
Promptagator 13.03 24.50 42.88 36.65 72.23 16.82 40.19
Document40% reg 12.54 24.94 44.89 $\uparrow$ 2.01 36.86 72.94 $\uparrow$ 0.71 18.16 $\uparrow$ 1.34 41.65 $\uparrow$ 1.46
Queryreg $\circ$ Document40% reg 14.66 $\uparrow$ 1.63 27.58 $\uparrow$ 3.08 47.41 $\uparrow$ 4.53 38.54 $\uparrow$ 1.89 74.04 $\uparrow$ 1.81 18.81 $\uparrow$ 1.99 42.45 $\uparrow$ 2.26
Document60% reg 14.11 $\uparrow$ 1.08 25.80 $\uparrow$ 1.30 43.99 $\uparrow$ 1.11 37.03 72.94 $\uparrow$ 0.71 16.84 41.77 $\uparrow$ 1.58
Queryreg $\circ$ Document60% reg 14.75 $\uparrow$ 1.72 28.34 $\uparrow$ 3.84 47.25 $\uparrow$ 4.37 39.43 $\uparrow$ 2.78 74.61 $\uparrow$ 2.38 20.48 $\uparrow$ 3.66 43.33 $\uparrow$ 3.14
Document80% reg 14.47 $\uparrow$ 1.44 25.53 $\uparrow$ 1.03 42.47 37.57 $\uparrow$ 0.92 72.61 $\uparrow$ 0.38 15.10 $\downarrow$ -1.72 41.55 $\uparrow$ 1.36
Queryreg $\circ$ Document80% reg 15.84 $\uparrow$ 2.81 27.96 $\uparrow$ 3.46 46.96 $\uparrow$ 4.08 39.09 $\uparrow$ 2.44 74.71 $\uparrow$ 2.48 20.49 $\uparrow$ 3.67 43.41 $\uparrow$ 3.22
Instrreg 13.80 $\uparrow$ 0.77 25.00 45.78 $\uparrow$ 2.90 36.29 73.53 $\uparrow$ 1.30 18.24 $\uparrow$ 1.42 41.87 $\uparrow$ 1.68
Queryreg $\circ$ Instrreg 15.20 $\uparrow$ 2.17 26.88 $\uparrow$ 2.38 47.44 $\uparrow$ 4.56 39.38 $\uparrow$ 2.73 74.08 $\uparrow$ 1.85 19.31 $\uparrow$ 2.49 42.52 $\uparrow$ 2.33
SPECTER-v2
Pretrained 13.59 25.63 41.92 35.91 71.46 21.74 38.84
Contrastive Fine-tuning 14.56 25.18 45.64 36.76 71.46 20.49 40.45
Promptagator 14.13 23.67 39.22 36.70 71.55 22.37 39.10
Document40% reg 15.39 $\uparrow$ 1.26 25.15 $\uparrow$ 1.48 43.39 $\uparrow$ 4.17 36.89 72.11 $\uparrow$ 0.56 21.66 $\downarrow$ -0.71 40.48 $\uparrow$ 1.38
Queryreg $\circ$ Document40% reg 14.19 24.99 $\uparrow$ 1.32 42.66 $\uparrow$ 3.44 36.39 71.49 21.40 $\downarrow$ -0.97 39.77 $\uparrow$ 0.67
Document60% reg 15.41 $\uparrow$ 1.28 26.79 $\uparrow$ 3.12 44.19 $\uparrow$ 4.97 36.90 72.81 $\uparrow$ 1.26 21.35 $\downarrow$ -1.02 41.07 $\uparrow$ 1.97
Queryreg $\circ$ Document60% reg 14.95 $\uparrow$ 0.82 25.81 $\uparrow$ 2.14 43.30 $\uparrow$ 4.08 38.38 $\uparrow$ 1.68 72.19 $\uparrow$ 0.64 21.60 $\downarrow$ -0.77 40.55 $\uparrow$ 1.45
Document80% reg 17.02 $\uparrow$ 2.89 28.63 $\uparrow$ 4.96 44.97 $\uparrow$ 5.75 38.35 $\uparrow$ 1.65 73.94 $\uparrow$ 2.39 22.78 $\uparrow$ 0.41 42.37 $\uparrow$ 3.27
Queryreg $\circ$ Document80% reg 14.70 $\uparrow$ 0.57 26.85 $\uparrow$ 3.18 44.66 $\uparrow$ 5.44 37.61 $\uparrow$ 0.91 73.07 $\uparrow$ 1.52 22.65 41.70 $\uparrow$ 2.60
Instrreg 14.73 $\uparrow$ 0.60 25.82 $\uparrow$ 2.15 41.50 $\uparrow$ 2.28 35.91 $\downarrow$ -0.79 73.06 $\uparrow$ 1.51 21.05 $\downarrow$ -1.32 39.24 $\uparrow$ 0.14
Queryreg $\circ$ Instrreg 14.50 $\uparrow$ 0.37 25.32 $\uparrow$ 1.65 41.44 $\uparrow$ 2.22 36.01 $\downarrow$ -0.69 72.46 $\uparrow$ 0.91 22.26 39.19
Table 4: DORIS-MAE Full Experimental Results
Model (Method) R@5 R@10 R@20 RP NDCG@10 MRR@10 MAP
E5-Large-v2
Pretrained 59.37 76.74 88.16 23.33 47.75 38.96 39.84
Contrastive Fine-tuning 70.69 85.86 94.77 29.34 56.15 47.13 47.57
Promptagator 69.41 83.35 94.11 30.38 55.14 46.67 47.24
Document40% reg 78.72 $\uparrow$ 9.31 90.43 $\uparrow$ 7.08 96.46 $\uparrow$ 2.35 39.40 $\uparrow$ 9.02 64.24 $\uparrow$ 9.10 56.45 $\uparrow$ 9.78 56.46 $\uparrow$ 9.22
Queryreg $\circ$ Document40% reg 77.76 $\uparrow$ 8.35 88.89 $\uparrow$ 5.54 96.21 $\uparrow$ 2.10 38.20 $\uparrow$ 7.82 63.00 $\uparrow$ 7.86 55.24 $\uparrow$ 8.57 55.38 $\uparrow$ 8.14
Document60% reg 78.16 $\uparrow$ 8.75 90.85 $\uparrow$ 7.50 96.41 $\uparrow$ 2.30 39.63 $\uparrow$ 9.25 64.41 $\uparrow$ 9.27 56.57 $\uparrow$ 9.90 56.53 $\uparrow$ 9.29
Queryreg $\circ$ Document60% reg 78.34 $\uparrow$ 8.93 90.52 $\uparrow$ 7.17 96.37 $\uparrow$ 2.26 39.06 $\uparrow$ 8.68 64.10 $\uparrow$ 8.96 56.24 $\uparrow$ 9.57 56.24 $\uparrow$ 9.00
Document80% reg 77.37 $\uparrow$ 7.96 89.81 $\uparrow$ 6.46 95.57 $\uparrow$ 1.46 38.14 $\uparrow$ 7.76 63.14 $\uparrow$ 8.00 55.16 $\uparrow$ 8.49 55.21 $\uparrow$ 7.97
Queryreg $\circ$ Document80% reg 78.44 $\uparrow$ 9.03 90.82 $\uparrow$ 7.47 96.25 $\uparrow$ 2.14 39.69 $\uparrow$ 9.31 64.36 $\uparrow$ 9.22 56.49 $\uparrow$ 9.82 56.45 $\uparrow$ 9.21
Instrreg 72.06 $\uparrow$ 2.65 86.23 $\uparrow$ 2.88 95.70 $\uparrow$ 1.59 32.14 $\uparrow$ 1.76 57.76 $\uparrow$ 2.62 49.17 $\uparrow$ 2.50 49.58 $\uparrow$ 2.34
Queryreg $\circ$ Instrreg 71.59 $\uparrow$ 2.18 85.12 $\uparrow$ 1.77 95.18 $\uparrow$ 1.07 31.65 $\uparrow$ 1.27 56.89 $\uparrow$ 1.75 48.37 $\uparrow$ 1.70 48.86 $\uparrow$ 1.62
SimCSE
Pretrained 49.03 68.92 81.12 15.65 39.23 30.35 31.42
Contrastive Fine-tuning 36.70 51.22 64.79 11.18 28.85 22.12 23.57
Promptagator 69.49 84.45 92.54 31.91 56.38 47.97 48.36
Document40% reg 74.92 $\uparrow$ 5.43 86.37 $\uparrow$ 1.92 93.36 $\uparrow$ 0.82 35.54 $\uparrow$ 3.63 60.24 $\uparrow$ 3.86 52.42 $\uparrow$ 4.45 52.60 $\uparrow$ 4.24
Queryreg $\circ$ Document40% reg 76.07 $\uparrow$ 6.58 87.46 $\uparrow$ 3.01 93.97 $\uparrow$ 1.43 36.96 $\uparrow$ 5.05 61.38 $\uparrow$ 5.00 53.57 $\uparrow$ 5.60 53.73 $\uparrow$ 5.37
Document60% reg 73.20 $\uparrow$ 3.71 84.71 92.10 33.47 $\uparrow$ 1.56 58.46 $\uparrow$ 2.08 50.58 $\uparrow$ 2.61 50.84 $\uparrow$ 2.48
Queryreg $\circ$ Document60% reg 75.75 $\uparrow$ 6.26 86.58 $\uparrow$ 2.13 93.19 $\uparrow$ 0.65 36.68 $\uparrow$ 4.77 60.94 $\uparrow$ 4.56 53.24 $\uparrow$ 5.27 53.44 $\uparrow$ 5.08
Document80% reg 70.38 81.60 $\downarrow$ -2.85 89.67 $\downarrow$ -2.87 31.42 55.76 47.97 48.39
Queryreg $\circ$ Document80% reg 74.10 $\uparrow$ 4.61 85.13 92.59 35.71 $\uparrow$ 3.80 59.79 $\uparrow$ 3.41 52.18 $\uparrow$ 4.21 52.45 $\uparrow$ 4.09
Instrreg 70.97 86.14 $\uparrow$ 1.69 93.91 $\uparrow$ 1.37 32.18 57.41 48.80 49.14
Queryreg $\circ$ Instrreg 71.46 $\uparrow$ 1.97 86.34 $\uparrow$ 1.89 93.78 $\uparrow$ 1.24 33.18 $\uparrow$ 1.27 58.16 $\uparrow$ 1.78 49.78 $\uparrow$ 1.81 50.04 $\uparrow$ 1.68
RoBERTa
Pretrained 30.51 40.56 49.68 9.84 23.56 18.49 19.54
Contrastive Fine-tuning 59.12 74.36 84.67 24.33 47.57 39.65 40.29
Promptagator 61.51 76.67 88.19 26.16 49.80 41.85 42.51
Document40% reg 67.38 $\uparrow$ 5.87 81.87 $\uparrow$ 5.20 91.03 $\uparrow$ 2.84 30.26 $\uparrow$ 4.10 54.54 $\uparrow$ 4.74 46.43 $\uparrow$ 4.58 46.85 $\uparrow$ 4.34
Queryreg $\circ$ Document40% reg 67.29 $\uparrow$ 5.78 81.64 $\uparrow$ 4.97 91.08 $\uparrow$ 2.89 29.21 $\uparrow$ 3.05 54.04 $\uparrow$ 4.24 45.85 $\uparrow$ 4.00 46.27 $\uparrow$ 3.76
Document60% reg 67.51 $\uparrow$ 6.00 81.78 $\uparrow$ 5.11 90.77 $\uparrow$ 2.58 30.50 $\uparrow$ 4.34 54.85 $\uparrow$ 5.05 46.85 $\uparrow$ 5.00 47.26 $\uparrow$ 4.75
Queryreg $\circ$ Document60% reg 67.71 $\uparrow$ 6.20 81.71 $\uparrow$ 5.04 90.99 $\uparrow$ 2.80 29.74 $\uparrow$ 3.58 54.41 $\uparrow$ 4.61 46.34 $\uparrow$ 4.49 46.73 $\uparrow$ 4.22
Document80% reg 67.82 $\uparrow$ 6.31 80.97 $\uparrow$ 4.30 90.18 $\uparrow$ 1.99 30.47 $\uparrow$ 4.31 54.62 $\uparrow$ 4.82 46.79 $\uparrow$ 4.94 47.20 $\uparrow$ 4.69
Queryreg $\circ$ Document80% reg 67.38 $\uparrow$ 5.87 81.42 $\uparrow$ 4.75 90.70 $\uparrow$ 2.51 30.09 $\uparrow$ 3.93 54.53 $\uparrow$ 4.73 46.59 $\uparrow$ 4.74 46.97 $\uparrow$ 4.46
Instrreg 64.47 $\uparrow$ 2.96 79.29 $\uparrow$ 2.62 90.09 $\uparrow$ 1.90 26.77 $\uparrow$ 0.61 51.49 $\uparrow$ 1.69 43.20 $\uparrow$ 1.35 43.78 $\uparrow$ 1.27
Queryreg $\circ$ Instrreg 65.14 $\uparrow$ 3.63 80.38 $\uparrow$ 3.71 90.26 $\uparrow$ 2.07 27.59 $\uparrow$ 1.43 52.25 $\uparrow$ 2.45 43.89 $\uparrow$ 2.04 44.39 $\uparrow$ 1.88
SPECTER-v2
Pretrained 38.91 56.57 71.79 11.06 30.91 23.23 24.65
Contrastive Fine-tuning 41.90 59.65 75.70 12.25 33.20 25.26 26.75
Promptagator 38.64 58.34 75.15 9.59 30.24 21.77 23.40
Document40% reg 56.09 $\uparrow$ 17.45 74.12 $\uparrow$ 15.78 86.32 $\uparrow$ 11.17 21.02 $\uparrow$ 11.43 45.00 $\uparrow$ 14.76 36.30 $\uparrow$ 14.53 37.18 $\uparrow$ 13.78
Queryreg $\circ$ Document40% reg 51.28 $\uparrow$ 12.64 69.75 $\uparrow$ 11.41 83.33 $\uparrow$ 8.18 17.57 $\uparrow$ 7.98 40.59 $\uparrow$ 10.35 31.81 $\uparrow$ 10.04 32.96 $\uparrow$ 9.56
Document60% reg 58.87 $\uparrow$ 20.23 75.45 $\uparrow$ 17.11 86.93 $\uparrow$ 11.78 22.11 $\uparrow$ 12.52 46.55 $\uparrow$ 16.31 37.88 $\uparrow$ 16.11 38.69 $\uparrow$ 15.29
Queryreg $\circ$ Document60% reg 54.22 $\uparrow$ 15.58 71.81 $\uparrow$ 13.47 84.97 $\uparrow$ 9.82 19.35 $\uparrow$ 9.76 42.90 $\uparrow$ 12.66 34.21 $\uparrow$ 12.44 35.26 $\uparrow$ 11.86
Document80% reg 61.02 $\uparrow$ 22.38 76.84 $\uparrow$ 18.50 87.48 $\uparrow$ 12.33 24.39 $\uparrow$ 14.80 48.68 $\uparrow$ 18.44 40.24 $\uparrow$ 18.47 40.94 $\uparrow$ 17.54
Queryreg $\circ$ Document80% reg 57.55 $\uparrow$ 18.91 74.67 $\uparrow$ 16.33 86.65 $\uparrow$ 11.50 22.00 $\uparrow$ 12.41 46.01 $\uparrow$ 15.77 37.44 $\uparrow$ 15.67 38.27 $\uparrow$ 14.87
Instrreg 43.26 $\uparrow$ 4.62 63.89 $\uparrow$ 5.55 80.16 $\uparrow$ 5.01 12.95 $\uparrow$ 3.36 34.80 $\uparrow$ 4.56 26.07 $\uparrow$ 4.30 27.50 $\uparrow$ 4.10
Queryreg $\circ$ Instrreg 38.57 60.47 $\uparrow$ 2.13 77.12 $\uparrow$ 1.97 10.72 $\uparrow$ 1.13 31.72 $\uparrow$ 1.48 23.15 $\uparrow$ 1.38 24.69 $\uparrow$ 1.29
Table 5: ArguAna Full Experimental Results
Model (Method) R@5 R@10 R@20 RP NDCG@10 MRR@10 MAP
E5-Large-v2
Pretrained 18.41 22.21 27.06 9.34 15.25 13.07 13.84
Contrastive Fine-tuning 37.25 43.79 49.85 24.29 33.16 29.87 30.74
Promptagator 41.53 48.34 54.66 27.23 37.07 33.56 34.43
Document40% reg 41.40 47.56 $\downarrow$ -0.78 53.69 $\downarrow$ -0.97 27.83 $\uparrow$ 0.60 37.01 33.72 34.56
Queryreg $\circ$ Document40% reg 45.50 $\uparrow$ 3.97 51.65 $\uparrow$ 3.31 58.08 $\uparrow$ 3.42 30.58 $\uparrow$ 3.35 40.50 $\uparrow$ 3.43 37.01 $\uparrow$ 3.45 37.86 $\uparrow$ 3.43
Document60% reg 44.15 $\uparrow$ 2.62 50.84 $\uparrow$ 2.50 55.69 $\uparrow$ 1.03 30.07 $\uparrow$ 2.84 39.64 $\uparrow$ 2.57 36.16 $\uparrow$ 2.60 36.93 $\uparrow$ 2.50
Queryreg $\circ$ Document60% reg 46.23 $\uparrow$ 4.70 52.61 $\uparrow$ 4.27 59.76 $\uparrow$ 5.10 31.75 $\uparrow$ 4.52 41.58 $\uparrow$ 4.51 38.13 $\uparrow$ 4.57 39.02 $\uparrow$ 4.59
Document80% reg 28.30 $\downarrow$ -13.23 33.42 $\downarrow$ -14.92 39.85 $\downarrow$ -14.81 16.04 $\downarrow$ -11.19 24.05 $\downarrow$ -13.02 21.13 $\downarrow$ -12.43 22.07 $\downarrow$ -12.36
Queryreg $\circ$ Document80% reg 44.80 $\uparrow$ 3.27 51.04 $\uparrow$ 2.70 58.21 $\uparrow$ 3.55 30.87 $\uparrow$ 3.64 40.29 $\uparrow$ 3.22 36.94 $\uparrow$ 3.38 37.85 $\uparrow$ 3.42
Instrreg 42.15 $\uparrow$ 0.62 48.33 53.02 $\downarrow$ -1.64 27.57 37.33 33.89 $\uparrow$ 0.33 34.67
Queryreg $\circ$ Instrreg 44.90 $\uparrow$ 3.37 50.38 $\uparrow$ 2.04 56.28 $\uparrow$ 1.62 29.07 $\uparrow$ 1.84 39.29 $\uparrow$ 2.22 35.79 $\uparrow$ 2.23 36.66 $\uparrow$ 2.23
SimCSE
Pretrained 16.96 20.76 24.98 9.62 14.58 12.67 13.44
Contrastive Fine-tuning 26.78 31.74 37.15 15.73 23.11 20.42 21.28
Promptagator 30.97 37.35 44.26 18.57 27.15 23.99 24.95
Document40% reg 32.10 $\uparrow$ 1.13 37.43 43.07 $\downarrow$ -1.19 20.29 $\uparrow$ 1.72 28.30 $\uparrow$ 1.15 25.45 $\uparrow$ 1.46 26.34 $\uparrow$ 1.39
Queryreg $\circ$ Document40% reg 34.63 $\uparrow$ 3.66 40.71 $\uparrow$ 3.36 47.44 $\uparrow$ 3.18 22.03 $\uparrow$ 3.46 30.64 $\uparrow$ 3.49 27.51 $\uparrow$ 3.52 28.46 $\uparrow$ 3.51
Document60% reg 33.03 $\uparrow$ 2.06 38.09 43.96 20.47 $\uparrow$ 1.90 28.87 $\uparrow$ 1.72 25.97 $\uparrow$ 1.98 26.89 $\uparrow$ 1.94
Queryreg $\circ$ Document60% reg 34.84 $\uparrow$ 3.87 40.84 $\uparrow$ 3.49 47.08 $\uparrow$ 2.82 22.16 $\uparrow$ 3.59 30.77 $\uparrow$ 3.62 27.64 $\uparrow$ 3.65 28.58 $\uparrow$ 3.63
Document80% reg 19.75 $\downarrow$ -11.22 24.46 $\downarrow$ -12.89 29.89 $\downarrow$ -14.37 10.20 $\downarrow$ -8.37 16.73 $\downarrow$ -10.42 14.33 $\downarrow$ -9.66 15.20 $\downarrow$ -9.75
Queryreg $\circ$ Document80% reg 33.76 $\uparrow$ 2.79 39.17 $\uparrow$ 1.82 45.63 $\uparrow$ 1.37 20.71 $\uparrow$ 2.14 29.34 $\uparrow$ 2.19 26.26 $\uparrow$ 2.27 27.21 $\uparrow$ 2.26
Instrreg 33.81 $\uparrow$ 2.84 39.23 $\uparrow$ 1.88 45.79 $\uparrow$ 1.53 21.61 $\uparrow$ 3.04 29.74 $\uparrow$ 2.59 26.79 $\uparrow$ 2.80 27.72 $\uparrow$ 2.77
Queryreg $\circ$ Instrreg 37.34 $\uparrow$ 6.37 43.37 $\uparrow$ 6.02 50.45 $\uparrow$ 6.19 24.57 $\uparrow$ 6.00 33.18 $\uparrow$ 6.03 30.02 $\uparrow$ 6.03 30.99 $\uparrow$ 6.04
RoBERTa
Pretrained 2.84 3.60 5.33 1.11 2.17 1.73 2.10
Contrastive Fine-tuning 23.59 29.09 35.27 13.80 20.69 18.10 19.05
Promptagator 27.23 33.07 39.14 15.83 23.69 20.78 21.74
Document40% reg 28.30 $\uparrow$ 1.07 34.47 $\uparrow$ 1.40 40.89 $\uparrow$ 1.75 16.94 $\uparrow$ 1.11 24.79 $\uparrow$ 1.10 21.81 $\uparrow$ 1.03 22.77 $\uparrow$ 1.03
Queryreg $\circ$ Document40% reg 32.83 $\uparrow$ 5.60 39.55 $\uparrow$ 6.48 46.38 $\uparrow$ 7.24 20.55 $\uparrow$ 4.72 29.05 $\uparrow$ 5.36 25.82 $\uparrow$ 5.04 26.79 $\uparrow$ 5.05
Document60% reg 28.67 $\uparrow$ 1.44 35.14 $\uparrow$ 2.07 41.49 $\uparrow$ 2.35 17.00 $\uparrow$ 1.17 25.14 $\uparrow$ 1.45 22.06 $\uparrow$ 1.28 23.03 $\uparrow$ 1.29
Queryreg $\circ$ Document60% reg 34.30 $\uparrow$ 7.07 40.86 $\uparrow$ 7.79 47.52 $\uparrow$ 8.38 21.29 $\uparrow$ 5.46 30.15 $\uparrow$ 6.46 26.83 $\uparrow$ 6.05 27.82 $\uparrow$ 6.08
Document80% reg 21.74 $\downarrow$ -5.49 27.23 $\downarrow$ -5.84 33.16 $\downarrow$ -5.98 12.11 $\downarrow$ -3.72 18.91 $\downarrow$ -4.78 16.35 $\downarrow$ -4.43 17.23 $\downarrow$ -4.51
Queryreg $\circ$ Document80% reg 33.69 $\uparrow$ 6.46 39.74 $\uparrow$ 6.67 46.23 $\uparrow$ 7.09 20.69 $\uparrow$ 4.86 29.41 $\uparrow$ 5.72 26.20 $\uparrow$ 5.42 27.18 $\uparrow$ 5.44
Instrreg 28.12 $\uparrow$ 0.89 34.36 $\uparrow$ 1.29 40.80 $\uparrow$ 1.66 16.28 $\uparrow$ 0.45 24.51 $\uparrow$ 0.82 21.46 $\uparrow$ 0.68 22.43 $\uparrow$ 0.69
Queryreg $\circ$ Instrreg 33.09 $\uparrow$ 5.86 39.65 $\uparrow$ 6.58 46.09 $\uparrow$ 6.95 20.57 $\uparrow$ 4.74 29.28 $\uparrow$ 5.59 26.06 $\uparrow$ 5.28 27.03 $\uparrow$ 5.29
SPECTER-v2
Pretrained 2.98 4.22 5.26 0.97 2.43 1.87 2.18
Contrastive Fine-tuning 7.85 9.60 12.56 3.72 6.36 5.35 5.89
Promptagator 6.56 8.46 11.81 3.60 5.72 4.88 5.40
Document40% reg 9.98 $\uparrow$ 3.42 12.52 $\uparrow$ 4.06 16.12 $\uparrow$ 4.31 4.99 $\uparrow$ 1.39 8.55 $\uparrow$ 2.83 7.32 $\uparrow$ 2.44 7.92 $\uparrow$ 2.52
Queryreg $\circ$ Document40% reg 10.84 $\uparrow$ 4.28 12.76 $\uparrow$ 4.30 16.26 $\uparrow$ 4.45 5.30 $\uparrow$ 1.70 8.81 $\uparrow$ 3.09 7.57 $\uparrow$ 2.69 8.18 $\uparrow$ 2.78
Document60% reg 8.64 $\uparrow$ 2.08 11.42 $\uparrow$ 2.96 14.49 $\uparrow$ 2.68 4.51 $\uparrow$ 0.91 7.46 $\uparrow$ 1.74 6.26 $\uparrow$ 1.38 6.80 $\uparrow$ 1.40
Queryreg $\circ$ Document60% reg 7.54 $\uparrow$ 0.98 10.09 $\uparrow$ 1.63 13.29 $\uparrow$ 1.48 4.03 $\uparrow$ 0.43 6.55 $\uparrow$ 0.83 5.48 $\uparrow$ 0.60 6.01 $\uparrow$ 0.61
Document80% reg 7.39 $\uparrow$ 0.83 10.27 $\uparrow$ 1.81 13.19 $\uparrow$ 1.38 3.62 6.46 $\uparrow$ 0.74 5.31 $\uparrow$ 0.43 5.87 $\uparrow$ 0.47
Queryreg $\circ$ Document80% reg 8.14 $\uparrow$ 1.58 11.28 $\uparrow$ 2.82 14.69 $\uparrow$ 2.88 4.43 $\uparrow$ 0.83 7.33 $\uparrow$ 1.61 6.13 $\uparrow$ 1.25 6.69 $\uparrow$ 1.29
Instrreg 10.49 $\uparrow$ 3.93 13.30 $\uparrow$ 4.84 16.80 $\uparrow$ 4.99 6.05 $\uparrow$ 2.45 9.30 $\uparrow$ 3.58 8.07 $\uparrow$ 3.19 8.67 $\uparrow$ 3.27
Queryreg $\circ$ Instrreg 11.08 $\uparrow$ 4.52 13.83 $\uparrow$ 5.37 17.36 $\uparrow$ 5.55 6.09 $\uparrow$ 2.49 9.59 $\uparrow$ 3.87 8.28 $\uparrow$ 3.40 8.91 $\uparrow$ 3.51
Table 6: WhatsThatBook Full Experimental Results